### v1

# Machine-generated text detection prevents language model collapse

Abstract
As Large Language Models (LLMs) become increasingly prevalent, their generated outputs are proliferating across the web, risking a future where machine-generated content dilutes human-authored text. Since web data is the primary resource for LLM pretraining, future models will be trained on an unknown portion of synthetic data. This will lead to model collapse, a degenerative process which causes models to reinforce their own errors and experience a drop in model performance. In this study, we investigate the impact of decoding strategy on model collapse, where we analyse the characteristics of the generated data during recursive training, its similarity to human references and the resulting model performance. Using the decoding strategies that lead to the most significant model degradation, we tackle the question: how to avoid model collapse when the origin (human or synthetic) of the training data is unknown. We design a novel methodology based on resampling the data distribution using importance weights from our machine-generated text detector. Our method is validated on two LLM variants (GPT-2 and SmolLM2) on the open-ended text generation task, demonstrating that we can successfully prevent model collapse and when there is enough human-authored data in the training dataset, our method improves model performance.
Machine-generated text detection prevents language model collapse
George Drayson and Vasileios Lampos Centre for Artificial Intelligence Department of Computer Science University College London, UK {george.drayson.23, v.lampos}@ucl.ac.uk
1 Introduction
Large Language Models (LLMs) can generate high quality, human-like language across a wide range of applications. A key factor that drives their capabilities is the vast amount of data used to train them, which is predominantly based on text published on the web Wenzek et al. (2020). The extensive adoption of LLMs will inevitably result in an ever-increasing amount of synthetic (machine-generated) data that will co-exist alongside or even dominate over human-generated text online Dohmatob et al. (2024b), especially within online ecosystems such as social media platforms, news websites, and online encyclopedias. Hence, there are legitimate concerns as to the effect this will have on future generations of language models trained on a mixed set of human and synthetic corpora.
Several works have attempted to simulate this scenario by recursively training language models on LLM-generated output Shumailov et al. (2023); Briesch et al. (2023); Alemohammad et al. (2024a). They show that this recursive training loop affects model performance, reflected by increasing perplexity among other metrics. The outcome of this recursive training process is referred to as “model collapse” Shumailov et al. (2023), a degenerative process caused by training on synthetic data from previous model generations, leading to models compounding errors over generations and converging to a distribution with very little variance and little resemblance to the original human data. The result is a model with significant performance degeneration Alemohammad et al. (2024a) and a drastic loss in diversity Briesch et al. (2023); Guo et al. (2024). Several works have explored how to mitigate model collapse, such as using synthetic samples as negative guidance in the image domain Alemohammad et al. (2024b), selecting synthetic samples based on high perplexity scores Feng et al. (2024) or filtering low-quality samples Zhang et al. (2024). However, in these prior works, the models are either trained on entirely synthetic data or the labels of the samples are known a priori.
In this work, we explore the scenario where the origin of the samples is unknown and we design a method of preventing model collapse by resampling the training data using an AI text detector. Our method is motivated by the results of Bertrand et al. (2024); Alemohammad et al. (2024a) which show that when the proportion of human data in the training set is high enough, model collapse can be prevented.
In addition to the current task-focused metrics used to assess model collapse, we evaluate the quality of the data from each model generation and its semantic resemblance to the original human samples. Using these metrics, we also explore the effect of decoding strategy on the extent of model collapse and the data characteristics that cause this. Our contributions can be summarised as follows:
  1. (a)
    we show that model collapse is significantly affected by the decoding strategy, demonstrating large discrepancies in performance and data quality,
  2. (b)
    we evaluate model collapse from three different perspectives: task performance, model generation quality, and semantic similarity to human text,
  3. (c)
    we train a machine-generated text detector to provide calibrated confidence estimates for the origin of text data,
  4. (d)
    we propose a method that incorporates the detector’s outputs to mitigate model collapse, and
  5. (e)
    present experiments on two LLMs and across a range of decoding strategies.
2 Related Work
2.1 Model collapse
Model collapse is a degenerative process in which models recursively trained on generational data exhibit a drop in performance compared to a model trained on the original human distribution Shumailov et al. (2023). In the early stages of recursive training, information is lost at the tails of the distribution and eventually, the data distribution converges to a point estimate with very little variance, resulting in a model that cannot be restored to the original model trained on the human data distribution. This effect can also be viewed as a change to neural scaling laws, in which there reaches a point where additional synthetic data does not improve model performance Dohmatob et al. (2024b).
It has been argued that the two causes for this behaviour are finite sampling error leading to information being lost at the tails of the distribution, and functional approximation error introducing non-zero likelihoods outside of the support of the original distribution Shumailov et al. (2023). Additionally, Dohmatob et al. (2024b) theorised that the choice of generation algorithm is another contributing factor to model collapse, however, this has not been empirically evaluated in the case of LLMs where decoding strategies that truncate the probability could have a significant impact on model collapse. Currently, model collapse in LLMs has been evaluated with a fixed decoding strategy and model degradation is assessed using task performance metrics including perplexity Shumailov et al. (2024) and test loss. An interesting perspective by Guo et al. (2024) is to evaluate the lexical and semantic diversity of LLMs after recursive training, presenting an interesting perspective to model collapse in LLMs - the data they generate. In this study we investigate model collapse in three dimensions: the generated data quality, its distributional similarity to human text and the downstream task performance of the model trained on this data.
2.2 Preventing Model collapse
With the increase in synthetically generated content on the internet, recent studies have explored methods of mitigating model collapse in generative models caused by training on synthetic data. Bertrand et al. (2024); Alemohammad et al. (2024a) show that when a high enough proportion of human data is added to the training data at each iteration, model collapse can be avoided in diffusion models. In the language domain Gerstgrasser et al. (2024) showed that by accumulating all cross-generational data and combining it with the original human data, model collapse can be significantly mitigated. Feng et al. (2024); Zhang et al. (2024) investigated model collapse in the purely synthetic setting, implementing methods of selecting synthetic samples based on heuristics with Feng et al. (2024) achieving performance akin to training on the original human data, avoiding model collapse.
However, in these prior works, the models are trained on either entirely synthetic data or the labels of the samples are known a priori. In this work, we investigate how to avoid model collapse in a more realistic setting where the training data is mixed and the origin (human or synthetic) of the samples and hence the class distribution is unknown.
3 Background
3.1 Language models and decoding strategies
Language models learn a probability distribution $p_{\theta}\!\left(\mathbf{x}\right)$ for a sequence of tokens $\mathbf{x}\!=\!\{x_{1},\dots,x_{|\mathbf{x}|}\}$, where $\theta$ denotes the model parameters. In this work, we study open-ended text generation, in which a sequence of $m$ tokens, $\mathbf{x}\!=\!\{x_{1},\dots,x_{m}\}$, is provided as context and the task is to generate a continuation, $\mathbf{\hat{x}}\!=\!\{\hat{x}_{1},\dots,\hat{x}_{c}\}$, from the model’s probability distribution, as in
$p_{\theta}\!\left(\hat{\mathbf{x}}\right)=\prod_{i=1}^{c}p_{\theta}\!\left(\hat{x}_{i}\mid\{\mathbf{x},\hat{\mathbf{x}}_{<i}\}\right)\,.$ (1)
Tokens are selected from the probability distribution at each step by following a decoding strategy, resulting in a text sample $\{\mathbf{x},\hat{\mathbf{x}}\}$.
There are two main categories of decoding strategies, deterministic and stochastic. The former is designed to maximise the joint probability of the generated sequence, e.g. by selecting the most probable token at each step (greedy decoding) or keeping track of multiple candidate text sequences and selecting the most probable (beam search). This reduces the chance of getting stuck in locally optimal but suboptimal sequences. Stochastic methods, on the other hand, produce less repetitive and more human-like text Holtzman et al. (2020). The simplest stochastic method samples directly from the distribution $p_{\theta}$. A more nuanced approach, referred to as top-$k$ decoding Fan et al. (2018), samples from the $k$ most probable tokens to avoid text generation from the tail of $p_{\theta}$. Nucleus sampling dynamically truncates the vocabulary to the highest probability tokens by thresholding the cumulative probability mass with a parameter $\eta\!\in\![0,1]$ Holtzman et al. (2020). Alternatively, the probability mass can be skewed towards high-probability events by deploying temperature, controlled by parameter $\tau\!\in\![0,1]$ Ackley et al. (1985).
3.2 Recursive LLM training
Similarly to Shumailov et al. (2024) and Dohmatob et al. (2024b), we simulate model collapse by training a language model recursively on its own generated output (entirely or partially, depending on our underlying hypothesis) for a fixed number of generations. This process is described in Algorithm 1. Recursive training commences by finetuning a pretrained language model, $p_{\theta}$, using a dataset consisting of $n$ human-generated samples, $\mathcal{D}_{\text{H}}=\{\mathbf{x}_{s}\}_{s=1}^{n}$. This results in a model $p^{0}$, where ‘00’ denotes the stage of the entire process (generation). [^1] For enhanced notational clarity, we choose to drop parameter $\theta$ for the recursively produced LLMs. However, we clarify that $\theta$ is updated in each generation. We then use a set of $n$ context sequences $\mathcal{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ (one for each sample in $\mathcal{D}_{\text{H}}$) to generate a set of continuation sequences $\mathcal{\hat{X}}=\{\mathbf{\hat{x}}_{1},\dots,\mathbf{\hat{x}}_{n}\}$ where $\hat{\mathbf{x}}_{s}\sim p^{0}$ (see also section 3.1). The human-generated context together with the LLM-generated continuation sequences form a new synthetic dataset, $\mathcal{D}_{\text{S}}^{1}$ (here ‘$1$’ is used to denote that this dataset will be used to finetune a language model in the next generation).
Subsequently, successive rounds of recursive training are carried out. In each generation $i$ with $i\!\geq\!1$, the original language model $p_{\theta}$ is finetuned using (an ever changing) synthetic dataset $\mathcal{D}_{\text{S}}^{i}$ to obtain $p^{i}$. Then, we prompt $p^{i}$ with context sequences $\mathcal{X}$ to generate a new synthetic dataset $\mathcal{D}_{\text{S}}^{i\!+\!1}$ that will be used to finetune $p_{\theta}$ in generation $i\!+\!1$.
Algorithm 1 Recursive LLM training
1:Input: Human text samples $\mathcal{D}_{\text{H}}=\{\mathbf{x}_{s}\}_{s=1}^{n}$, pretrained language model $p_{\theta}$
2:Obtain $p^{0}$ by finetuning $p_{\theta}$ using $\mathcal{D}_{\text{H}}$
3:for $i=1,\dots,G$ do
4:     $\mathcal{D}_{\text{S}}^{i}=\{\mathbf{x}_{s},\hat{\mathbf{x}}_{s}\}_{s=1}^{n},\text{ where }\hat{\mathbf{x}}_{s}\sim p^{i-1}$
5:     Obtain $p^{i}$ by finetuning $p_{\theta}$ using $\mathcal{D}_{\text{S}}^{i}$
6:end for
7:Outputs: $p^{i}\,(i\!\geq\!0)$, $\mathcal{D}_{\text{S}}^{i}\,(i\!\geq\!1)$
3.3 LLM performance
We take inspiration from  Shumailov et al. (2024); Dohmatob et al. (2024b) and evaluate model collapse by fine-tuning and evaluating models on the open-ended text generation task. We assess language modelling performance in terms of perplexity and evaluation accuracy. Perplexity measures how well the model predicts unseen text, with lower values indicating better predictive performance. Accuracy, in this context, reflects the proportion of correctly predicted tokens, providing a direct measure of the model’s effectiveness in generating accurate language outputs.
3.4 Qualitative metrics of generated text outputs
We complement performance metrics with more qualitative ones drawn on the generated text outputs and the similarity to human references. This helps us obtain a holistic understanding of model collapse.
Diversity (D) takes into account the sequence-level repetition at different $n$-gram levels of a document. Higher scores are reflective of more lexically diverse text. We use the following formulation
$\texttt{D}\!\left(\hat{\mathbf{x}}\right)=\prod_{n=2}^{4}\left(1-\frac{\lvert\text{unique }n\text{-grams}(\hat{\mathbf{x}})\rvert}{\lvert\text{total }n\text{-grams}(\hat{\mathbf{x}})\rvert}\right)\,.$ (2)
MAUVE Pillutla et al. (2021) measures the distribution similarity between the original human text and the generated text. It is computed using the Kullback–Leibler (KL) divergence between the two text distributions in the embedding space of a large language model. We calculate MAUVE using the implementation of Pillutla et al. (2021) over a random sample of $1{,}000$ documents of human and machine-generated text. A higher score indicates that the model generates more human-like text.
Self-BLEU Zhu et al. (2018) evaluates the BLEU score Papineni et al. (2002) of each document compared to all other documents in the generation set, providing a metric for how repetitive the model is for different outputs. Due to the computational expense of evaluating Self-BLEU we constrain the calculation to a random sample of $1{,}000$ documents and evaluate up to an $n$-gram size of $4$. A lower Self-BLEU score indicates higher text diversity.
Readability is evaluated using the Flesch-Kincaid Reading-Ease score Flesch (1948), which evaluates how difficult a passage is to understand based on the number of words, sentences, and syllables. We implement the metric using the textstat package  [^2] https://textstat.org/. Lower scores indicate more complex text that contains longer sentences and verbose language.
4 The impact of decoding strategies on model collapse
Figure 1: Perplexity and accuracy of GPT-2 (top) and SmolLM2 (bottom) over generations of fully synthetic recursive training.
Model Decoding Diversity $\uparrow$ S-BLEU $\downarrow$ MAUVE $\uparrow$ Readability $\uparrow$
GPT-2 greedy $0.96$/$0.70$ $61.02$/$67.13$ $0.99$/$1.00$ $60.47$/$8.25$
beam search $16.78$/$10.86$ $61.54$/$67.60$ $0.91$/$1.29$ $60.97$/$17.57$
sampling $94.88$/$99.82$ $24.12$/$6.76$ $90.16$/$7.18$ $40.62$/$-10.14$
temperature $87.76$/$25.10$ $33.45$/$54.56$ $94.15$/$22.69$ $46.80$/$\mathbf{36.79}$
top-$k$ $84.57$/$70.20$ $38.81$/$\mathbf{42.14}$ $95.21$/$\mathbf{70.01}$ $51.19$/$34.32$
nucleus $92.26$/$\mathbf{86.73}$ $28.24$/$26.86$ $90.96$/$57.41$ $43.96$/$21.31$
SmolLM2 greedy $6.99$/$3.02$ $61.90$/$50.23$ $4.29$/$1.12$ $62.33$/$\mathbf{46.72}$
beam search $6.75$/$2.88$ $61.47$/$\mathbf{44.47}$ $3.86$/$1.09$ $62.38$/$44.47$
sampling $90.83$/$\mathbf{91.62}$ $37.69$/$33.33$ $91.28$/$\mathbf{83.82}$ $47.49$/$44.33$
temperature $82.96$/$24.37$ $46.21$/$61.64$ $91.48$/$21.66$ $52.78$/$64.59$
top-$k$ $82.57$/$57.35$ $48.62$/$61.18$ $95.48$/$68.41$ $55.45$/$62.82$
nucleus $87.38$/$47.36$ $42.43$/$60.85$ $94.40$/$67.37$ $51.14$/$63.71$
Human $88.79$ $42.89$ $100$ $50.34$
Table 1: Impact of decoding strategies on the generation quality of GPT-2 and SmolLM2 from generation 00 to $9$ of fully synthetic recursive training. For each model we highlight in bold the closest score to the human reference for generation $9$ ($\uparrow$ / $\downarrow$: higher / lower is better).
We evaluate model collapse on the open-ended text generation task by finetuning LLMs on the Wiki- Text-2 dataset Merity et al. (2016) with GPT-2 and Hugging Face’s recently released SmolLM2 Allal et al. (2025). The Wikipedia articles are segmented into chunks of $512$ tokens where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full finetuning for 1 epoch and to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample Dohmatob et al. (2024b). Additional details about the experimental setup can be found in Appendix B.1.
As described in section 3.2, we recursively train language models on the previous generation’s output. We evaluate a range of decoding strategies to assess the effect on model collapse, including greedy decoding ($\tau=0$), $5$-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Garces Arias et al. (2025).
In Figure 1 we present the perplexity and evaluation accuracy on the WikiText test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1). We present the qualitative metrics at generation 0 and 9 in Table 1.
Deterministic decoding results in the most significant model collapse. While stochastic sampling appear to follow a linear relationship in model degradation across generations, the rate of collapse is accelerated with greedy decoding and beam search before plateauing in later generations. The data quality produced by the models provides some intuition for this as the starting metrics are significantly lower than stochastic methods with diversity and MAUVE scores less than $10\%$ for SmolLM2, evidence of repetitive and non-fluent language. Due to the poor data quality and exacerbated model collapse we choose not to carry out our subsequent experiments with deterministic strategies.
Pure sampling affected the models in opposite ways. In the case of GPT-2, sampling directly from the probability distribution leads to diverse and fluent text in generation 00 but the effect of training on this data is the highest perplexity among all stochastic methods and generated text that does not resemble anything similar to the human reference texts with a final MAUVE of $7.18$. With SmolLM2, on the other hand, pure sampling led to the smallest drop in model performance and the highest quality data on average across all metrics.
Recursive training with temperature leads to the most repetitive text of all stochastic methods with a ${}70\%$ decrease in diversity for both models and the furthest distribution semantically from the human reference for SmolLM2, showing significant model collapse.
Performance with top-$k$ was consistent across models, with the smallest drop in diversity and Self-BLEU, the closest resulting text to the human reference and a smaller drop in test performance compared to Nucleus sampling.
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most extreme scenario while also facilitating direct comparison between models. For this reason, we evaluate the models using the worst performing stochastic decoding method (pure sampling for GPT-2 and temperature sampling for SmolLM2) and top-$k$ decoding due to its consistent performance across models that will allow for direct comparison.
5 Machine-generated text detector
To prevent model collapse in the setting where the origin of each training sample is unknown, we train a detector to estimate the likelihood of the text being machine-generated. In the following section, we define how we trained our model, followed by our method of weighted sampling using the detector outputs as importance scores.
Machine-generated text detection methods can be divided into neural-based and metric-based approaches. Metric-based methods use statistical features, often extracted from surrogate LLMs, to detect machine-generated text, whereas neural-based methods are implemented by fine-tuning a small pre-trained language model with a binary classification head. We decided on a neural-based approach due to their state-of-the-art performance on common benchmarks Wang et al. (2024); Li et al. (2024) and the ability to use the model output logits as a form of uncertainty.
Model in-distribution out-of-distribution
AUC Acc. F1 AUC Acc. F1
ModernBERT 98.61 94.79 94.80 94.27 86.07 86.01
DeBERTaV3 97.10 95.43 95.42 81.68 81.18 80.96
RoBERTa 98.15 94.03 94.04 84.60 80.64 80.36
Table 2: Machine-generated text detection performance on MAGE. The F1-score and accuracy (Acc.) are calculated as macro-averages.
Figure 2: GPT2 (top) and SmolLM2 (bottom) under partially synthetic recursive training ($\alpha=1$) for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
[…]
Table 3: Comparison of data quality when using our mitigation strategy versus the baseline for GPT-2 and SmolLM2. Results are shown for top-$k$ decoding and pure sampling/temperature with different values for $\alpha$, $\beta$, and $\gamma$. The proportion of human data in the recursive training is given by $\alpha$. $\uparrow$ means higher is better and $\downarrow$ means lower is better. Positive improvements are shown in blue and negative results in red.
5.1 Model
Our detector is based on an encoder-only language model with a sequence classification head that maps the CLS token representation to logits, $\mathbf{z}_{i}$, which are converted to probabilities using a sigmoid function. As LLM training is considerably resource intensive, any data filtering or sampling methods must be able to efficiently process large quantities of data with minimal computational overhead Wenzek et al. (2020). Based on these requirements we evaluated the base variants of three model architectures with under 200 million parameters: RoBERTa Goyal et al. (2020) and DeBERTav3 He et al. (2023) due to their state-of-the-art performance in machine-generated text detection benchmarks and ModernBERT Warner et al. (2024), a very recent model that has achieved superior performance on a range of Natural Language Inference and Information Retrieval benchmarks. The added advantage of ModernBERT is the large context window (8192 tokens) and superior computational speed and memory efficiency Warner et al. (2024). More details on the models can be found in Appendix A.
Despite their strong performance, as with all modern neural networks, the confidence estimates are poorly calibrated, often resulting in overconfident predictions  Guo et al. (2017). To mitigate overconfidence in the predictions, we applied label smoothing to the target labels during training. Additionally, we implemented temperature scaling to further calibrate the model’s predictions. Given the logit vector $\mathbf{z}_{i}$, the new confidence prediction is $\hat{q}_{i}=\sigma_{\text{SM}}(\mathbf{z}_{i}/T)$ where $T$ is the temperature parameter optimised using negative log-likelihood loss on the validation set.
5.2 Training
Each model was trained and evaluated on the MAGE dataset Li et al. (2024), a popular AI text detection benchmark dataset based on 10 Natural Language Processing datasets which have been used to generate text from 27 LLMs. This dataset has already been split into train, validation and test splits in the ratio 80/10/10. We also test on their challenging out-of-distribution test set containing human-curated text from four unseen domains and machine-generated samples by an unseen model (GPT-4). Each model was finetuned for 5 epochs using a binary cross-entropy loss with early stopping. More details on the model training hyperparameters and the dataset can be found in Appendix A. The results of our evaluation can be seen in Table 2 where ModernBERT achieved the best classification performance on both test sets with an Area Under the Receiver Operating Characteristic Curve (AUC) of $98.61$ and $94.27$ respectively.
6 Preventing Model Collapse
So far we have carried out recursive training in the most extreme scenario where models are trained exclusively on the outputs of the previous generation and no human data is added during training. We now turn our focus to the partially synthetic setting, a more realistic scenario where human data make up a portion of the training dataset and the synthetic data is mixed across generations. In this setting, we explore model collapse with the assumption that the data distribution is unknown. The training dataset for generation $i$, $\mathcal{D}^{i}$, is defined in equation 3 with the three different scenarios depending on the mixing coefficients $\alpha$, $\beta$, and $\gamma\!\in\![0,1]$:
$\begin{split}\mathcal{D}^{i}\sim&\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{H}},\alpha\\ &\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{S}}^{i},\beta\\ &\,\,{\text{sample}_{i\geq 2}}\,\,\left\{\mathcal{D}_{\text{S}}^{i-1},\dots,\mathcal{D}_{\text{S}}^{1}\right\},\frac{\gamma}{(i-1)}\,.\end{split}$ (3)
  1. 1.
    Fully synthetic $\left(\alpha=0,\beta=1,\gamma=0\right)$: training data consists entirely of synthetic samples from the previous model generation.
  2. 2.
    Partially synthetic $\left(\alpha>0,\beta=1,\gamma=0\right)$: the same proportion of human data is added to the training data at every iteration.
  3. 3.
    Partially synthetic with cross-generational data $\left(\alpha>0,\beta>0,\gamma>0\right)$: as above but the synthetic data is accumulated across generations.
We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha,\beta,\gamma$ to affect the distribution of human and machine-generated data as well as the proportion of cross-generational data in the training set. We denote the sigmoid output from the detector for a sample $x_{n}$ as $q\left(x_{n}\right)$. Given a dataset $\mathcal{D}$, consisting of an unknown distribution of human and synthetic samples, and the predictions from the detector for each sample $q\left(x_{n}\right)\forall\mathcal{D}$, the goal of our method is to sample a new dataset which when used to finetune model $p_{\theta}$ does not result in model collapse. A model has collapsed if the addition of synthetic samples in the training data leads to degraded model performance compared to training exclusively on the human samples. We evaluate model performance and data quality as described in Sections 3.3 and 3.4.
6.1 Method
Our method is based on weighted sampling with replacement where the weights are extracted from the AI text detector model probabilities. As the detector has been trained on an unbalanced dataset ($71\%$ machine-generated), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$. To shift the probabilities in favour of predicting human samples, we apply a bias term (calculated based on the confidence threshold) to the probabilities, $b$, followed by normalising the weights so they sum to 1, as in
$w_{i}=\frac{\left(1-q\left(x_{n}\right)\right)^{b}}{\sum_{x_{j}\in D}\left(1-q\left(x_{j}\right)\right)^{b}}\quad\forall x_{n}\in D.$ (4)
Finally, $kD$ samples are drawn independently with replacement from $D$ using weights $w$, where $k$ was set to $1.5$ to allow for upsampling of the training data. To avoid oversampling specific samples we set the maximum number of selections to $10$.
6.2 Results
As explained in Section 4 we evaluate our method using the decoding strategy that led to the largest model collapse in each model, pure sampling and temperature sampling for GPT-2 and SmolLM2 respectively. In addition, we also implement the experiments with top-$k$ decoding to allow for direct comparison.
Our method is evaluated against the baseline of training on all samples in the pool of data, $\mathcal{D}$. In addition, we define the “Oracle” performance which represents a perfect AI text detector that filters all synthetic samples. The task performance metrics, data quality metrics and detector accuracy over generations of partially synthetic training $\left(\alpha=1,\beta=1,\gamma=0\right)$ can be seen in Figure 2. We also display the performance in the setting with half the human data $\alpha=0.5$ in Appendix S1. Over $9$ iterations of recursive training, our method of weighted sampling can be seen to prevent model collapse and when $\alpha=1$ i.e. there is an equal proportion of human to synthetic data, our method improves model performance compared to training on exclusively human data (Oracle) across both models and respective decoding strategies. In addition to the improved task performance, our method can also be seen to remove any effects on the readability, diversity and detectability of the synthetic samples over generations compared to the baseline. Improving performance on the Oracle demonstrates the value of synthetic data in LLM training. However, without our method resampling the distribution to contain more human-generated samples and removing likely synthetic text, the model collapses.
Figure 3: Perplexity Distribution for machine-generated text of GPT-2 at generations 0, 1 and 9 in the partially synthetic scenario ($\alpha=0.5$). The graphs on the left are the baseline and the graphs on the right are when our method is used. Perplexity is evaluated using the model trained on the human text. Machine-generated text from GPT-2 using top-$k$ decoding and sampling.
We evaluate recursive training under three different mixing coefficient settings. In Table 3 we present the percentage difference between the final model under the baseline strategy and using our method. From the table, it can be seen that mixing cross-generational data actually has an insignificant effect on the extent of model collapse compared to training with previous generational data only, contrasting with the findings of Gerstgrasser et al. (2024). However, in their experiments, the compute budget was unlimited whereas we take the more realistic setting of creating a fixed-size dataset using mixing coefficients.
We evaluate the perplexity distribution of the synthetic data at each generation using the model $p^{0}$ trained on the human data, similar to Shumailov et al. (2024). We plot in Figure 3 the distributions for generation 00, $1$ and $9$ for the GPT-2 model with top-$k$ and pure sampling and compare our method to the baseline. For top-$k$ decoding, we observe a similar behaviour to the results of Shumailov et al. (2024) in which the perplexity shifts towards regions of lower perplexity and the distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we see a completely different behaviour, the distributions shift right and become more spread out into regions of low probability and high perplexity. This is an interesting finding that demonstrates that by removing truncation from the decoding strategy, the narrowing effect of model collapse is removed and instead, we see the model collapse in a completely different way by generating long-tail incoherent text that is completely distinct from the original human samples. Using our mitigation strategy, however, we observe very little change in the perplexity distribution for both sampling strategies and improved data quality metrics to the baseline, as seen in Table 3.
7 Conclusion
In this work we investigate model collapse across three dimensions: model performance, data quality and resemblance to human samples. Through our analysis we found that the extent of model collapse and the effect on the data distribution is influenced by the decoding strategy where truncating the distribution can lead to peaked distributions and repetitive models while sampling directly can result in high perplexity and verbose outputs with low resemblance to human data. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario where human data is mixed into the training data, but in an unknown proportion. We designed a novel method to mitigate model collapse based on resampling the training distribution using the predictions of a machine-generated text detector. We have validated our method on two popular LLMs (GPT-2 and SmolLM2) across a range of decoding strategies, showing that we can prevent model collapse in all cases. When the proportion of human data is high enough, our method results in improved model performance compared to training only on the human data.
Limitations
The primary limitation of this study is the evaluation of relatively small LLMs (137M and 360M parameters) compared to larger models, which often have billions of parameters, limiting the generalisability of these findings. However, recent work Dohmatob et al. (2024a) suggests that model collapse is even more pronounced in larger models, indicating that our proposed approach could be even more beneficial in such cases. Future work could extend this analysis to larger models to validate these findings.
Additionally, similar to previous studies Shumailov et al. (2023); Dohmatob et al. (2024b), we assess LLMs exclusively in a fine-tuning setting rather than pretraining from scratch. While pretraining experiments could provide deeper insights, the computational cost and complexity of training large-scale models from the ground up make such an approach impractical in our case. Nevertheless, given that model collapse has been primarily evaluated in LLMs from a finetune setting, the conclusions made in this work still align with the current body of research.
Another limitation is that our study focuses primarily on open-ended text generation tasks. While this is a crucial area for understanding model collapse, our findings may not fully generalise to other domains, such as structured prediction or code generation, where the impact of model collapse may manifest differently. Future work could explore whether our resampling method remains effective across these domains.
Ethics Statement
The authors declare no competing interests. The Wikitext-2 dataset Merity et al. (2016) used for the experiments is licensed under the Creative Commons Attribution-ShareAlike License (CC BY-SA 4.0). GPT-2 is licensed under the MIT license and SmolLM2 is licensed under APACHE 2.0. AI assistants were not used for this research.
Appendix
Appendix A AI text detection
A.1 Pretrained models
RoBERTa
(Robustly Optimized BERT Pretraining Approach) Goyal et al. (2020) improves on the pre-training phase of BERT Kenton and Toutanova (2019), an encoder-only Transformer model that leverages masked language models to enable pre-trained deep bidirectional representations. The RoBERTa model optimised the pre-training procedure for BERT by training the model for longer and on more data, changing the masking pattern, and removing the next sentence prediction objective. We use the RoBERTa base variant which has $125$ million parameters.
DeBERTav3
(Decoding-enhanced BERT with Disentangled Attention) He et al. (2023), is a BERT-based encoder only model enhanced with disentangled attention. DeBERTav3 model improves on DeBERTa by using Enhanced Mask Decoding and an ELECTRA-style pretraining objective, Replaced Token Detection, instead of Masked Language Modeling. We use the base variant which contains $86$ million backbone parameters with an embedding layer of $98$ million parameters.
ModernBERT
Warner et al. (2024), is a recent addition to the encoder-only transformer models that has been designed to increase downstream performance and efficiency on GPUs, particularly in long-context scenarios due to its 8192 native sequence length. The model was trained on 2 trillion tokens and improves on the original BERT architecture with rotary positional embeddings (RoPE), unpadding, GeGLU layers and alternating local-global attention demonstrating state-of-the-art performance amongst encoder models across a range of classification and retrieval tasks. We conduct experiments with the base variant, which contains $150$ million parameters.
A.2 Dataset
The data used for the human text consists of opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD) and Scientific writing (SciGen). They sample 1000 texts from each domain (apart from opinion statements and news articles with 804 and 777 samples respectively) and generate text from 27 LLMs from 7 model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience and EleutherAI. For each human-written sample in the dataset, they generate a machine-generated version by providing the LLM the first 30 tokens of human-written text. In addition, for the OpenAI models, they investigate two other prompt strategies for applicable domains. These include ’topical’ prompts such as an argument or news title and ’specified’ prompts which contain information about the domain source. This results in 33,000 (27,000+3*2*1000) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio $80:10:10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source so that both sets have a roughly 50:50 balance. The training set, however, is $71\%$ machine-generated. All splits combined amount to $154,078$ human-written and $294,381$ machine-generated texts. In addition to the test set, we also evaluate our detector on their challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) and generated by an unseen model (GPT4).
[…]
GPT-2 Greedy 29.29/82.74 38.72/34.93 0.96/0.70 61.02/67.13 0.99/1.00 60.47/8.25
Beam Search 29.25/78.06 38.75/35.21 16.78/10.86 61.54/67.60 0.91/1.29 60.97/17.57
Sampling 29.29/58.64 38.74/32.49 94.88/99.82 24.12/6.76 90.16/7.18 40.62/-10.14
Temperature 29.23/44.55 38.77/34.47 87.76/25.10 33.45/54.56 94.15/22.69 46.80/36.79
Top-$k$ 29.31/48.36 38.73/33.12 84.57/70.20 38.81/42.14 95.21/70.01 51.19/34.32
Nucleus 29.28/48.96 38.74/32.73 92.26/86.73 28.24/26.86 90.96/57.41 43.96/21.31
SmolLM2 Greedy 13.36/90.24 47.89/44.00 6.99/3.02 61.90/50.23 4.29/1.12 62.33/46.72
Beam Search 13.36/88.90 47.87/43.96 6.75/2.88 61.47/44.47 3.86/1.09 62.38/44.47
Sampling 13.36/14.76 47.88/46.82 90.83/91.62 37.69/33.33 91.28/83.82 47.49/44.33
Temperature 13.36/23.70 47.86/46.38 82.96/24.37 46.21/61.64 91.48/21.66 52.78/64.59
Top-$k$ 13.36/19.01 47.88/46.41 82.57/57.35 48.62/61.18 95.48/68.41 55.45/62.82
Nucleus 13.36/21.80 47.88/46.55 87.38/47.36 42.43/60.85 94.40/67.37 51.14/63.71
Human 88.79 42.89 100 50.34
Table S1: Performance results and qualitative scores for GPT-2 and SmolLM2 over a range of decoding strategies. Generation 0 and 9 are presented and recursive training is in the fully synthetic scenario.
A.3 Hyperparameters
Each model was finetuned for 5 epochs with an early stopping patience of 3 epochs which is based on the macro average F1 score on the validation set and with no warm up. Models were evaluated on the validation set after every epoch which was used to select the best model for evaluation. Optimization was performed using AdamW with a $\beta_{1}$, $\beta_{2}$, $\epsilon$ and weight decay of $0.9$, $0.98$, $1e^{-6}$ and $1e^{-2}$ respectively. These parameters were chosen based on the recommended parameters of Warner et al. (2024). The label smoothing parameter $\alpha$ was set to 0.1. The seed was fixed at $42$ and a training batch size of $16$ and validation and test batch size of $64$ was used for all experiments. The learning rate was set based on a hyperparameter sweep over [$1e^{-5}$, $1.5e^{-5}$, $2e^{-5}$, $3e^{-5}$, $4e^{-5}$]. For ModernBERT the best performing learning rate was $1e^{-5}$. The classification threshold was selected to maximise the macro average F1 score on the validation set.
Appendix B Model collapse experiments
B.1 Model collapse training
In our experiments, we conduct full fine-tuning on the GPT-2 base (128 million parameters) and SmolLM2 (360 million parameters) using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we use $\beta_{1}$, $\beta_{2}$ and $\epsilon$ of $0.9$, $0.999$ and $081e-$ 08 respectively. Each model is trained for $1$ epoch with the hyperparameters fixed for all experiments. We conduct 10 iterations of recursive training.
B.2 Open-ended text generation dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2016). We train models on the training set, consisting of $36718$ documents and evaluate on the test set of $4358$ documents. The WikiText dataset was extracted from the Good or Featured article criteria specified by editors on Wikipedia and only covers the English language.
B.3 Models
We conduct experiments on the base variant of GPT-2 ($137$ million parameters) and the $360$ million parameter variant of SmolLM2.
Figure S1: GPT2 (top) and SmolLM2 (bottom) under partially synthetic recursive training (alpha=0.5) for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
[…]
GPT-2 Top-$k$ 0.5 1 0 84.46 / 82.16 39.10 / 39.55 93.31 / 91.10 51.30 / 46.16
0.5 0.5 0.5 85.66/84.35 38.94/40.64 94.07/92.00 51.06/45.38
1 1 0 84.66 / 82.68 39.19 / 40.20 93.24 / 92.69 50.99 / 48.09
Sampling 0.5 1 0 94.86 / 97.77 24.09 / 14.52 91.08 / 46.87 40.64 / 23.69
0.5 0.5 0.5 94.95/97.50 24.15/17.02 90.07/76.01 41.02/29.08
1 1 0 94.88 / 96.86 24.04 / 18.12 91.18 / 75.64 40.71 / 31.49
GPT-2 mitigation Top-$k$ 0.5 1 0 84.43 / 85.89 39.43 / 38.85 94.59 / 91.52 50.79 / 50.33
0.5 0.5 0.5 85.18/86.40 39.28/38.90 94.75/95.67 51.31/49.99
1 1 0 84.81 / 85.65 38.85 / 38.76 94.92 / 93.95 51.28 / 51.34
Sampling 0.5 1 0 94.87 / 96.50 24.30 / 19.85 92.92 / 81.58 40.68 / 35.59
0.5 0.5 0.5 94.92/96.46 24.09/20.66 91.16/88.46 40.98/36.48
1 1 0 94.61 / 96.17 24.01 / 21.87 91.76 / 88.06 40.78 / 38.10
SmolLM2 Top-$k$ 0.5 1 0 82.50/73.99 48.99/52.77 94.58/91.10 55.65/58.09
0.5 0.5 0.5 82.46/76.24 48.66/52.44 93.68/95.13 55.24/57.88
1 1 0 82.86/77.80 48.68/51.36 94.76/89.06 55.43/57.27
Temperature 0.5 1 0 83.42/66.25 45.57/52.77 95.37/81.27 52.72/57.67
0.5 0.5 0.5 83.27/71.22 45.89/52.73 88.33/84.54 53.01/56.23
1 1 0 82.73/72.75 45.69/52.21 94.57/87.55 52.88/56.39
SmolLM2mitigation Top-$k$ 0.5 1 0 82.21/81.77 49.04/49.91 94.96/93.19 55.49/56.64
0.5 0.5 0.5 82.56/81.86 48.83/49.62 94.55/94.46 55.33/56.55
1 1 0 82.35/82.52 45.00/48.91 94.00/94.43 55.35/56.14
Temperature 0.5 1 0 82.83/79.35 46.08/49.15 93.13/70.63 52.87/56.08
0.5 0.5 0.5 83.31/79.62 45.64/49.34 93.07/96.14 52.79/55.68
1 1 0 83.24/80.87 45.94/48.50 93.93/78.36 52.64/54.90
Table S2: Comparison of data quality when using our mitigation strategy versus the baseline. Results are shown for top-k decoding and pure sampling with different values for $\alpha$, $\beta$, and $\gamma$. The proportion of human data in the recursive training is given by $\alpha$. $\uparrow$ means higher is better and $\downarrow$ means lower is better.